OpenStudy (natsent06):

NEED HELP WITH QUESTIONS Derek’s favorite clothing store is having their annual fall sale. All clothing is 25 percent off. A shirt he really likes has an original price of $28. Derek has $24 in his wallet. He is wondering if he has enough money to buy the shirt now that it is on sale. He knows he must also pay sales tax. The sales tax rate is 8 percent and is calculated on the discounted price. Does Derek have enough money to buy the shirt? Justify your answer. Your explanation should contain answers to the following questions: 1. What is the amount of the discount? 2. What is the sale price? 3. How much wil

OpenStudy (anonymous):

Okay, divide 25 by 100 to get 0.25. Multiply that by 28, so we have: 28 * 0.25 = ??

OpenStudy (anonymous):

25% of 28 = 7, 28 - 7 = 21. 8% of 21 = 1.68, 1.68 + 21 = 22.68.

OpenStudy (anonymous):

I think that's right.